# SimPO: Simple Preference Optimization with a Reference-Free Reward

### v2

@@ Abstract
Direct Preference Optimization (DPO) is a widely used offline preference optimization algorithm that reparameterizes reward functions in reinforcement learning from human feedback (RLHF) to enhance simplicity and training stability. In this work, we propose SimPO, a simpler yet more effective approach. The effectiveness of SimPO is attributed to a key design: using the average log probability of a sequence as the implicit reward. This reward formulation better aligns with model generation and eliminates the need for a reference model, making it more compute and memory efficient. Additionally, we introduce a target reward margin to the Bradley-Terry objective to encourage a larger margin between the winning and losing responses, further enhancing the algorithm’s performance. We compare SimPO to DPO and its latest variants across various state-of-the-art training setups, including both base and instruction-tuned models like Mistral and Llama3. We evaluate on extensive instruction-following benchmarks, including AlpacaEval 2, MT-Bench, and the recent challenging Arena-Hard benchmark. Our results demonstrate that SimPO consistently and significantly outperforms existing approaches without substantially increasing response length. Specifically, SimPO outperforms DPO by up to 6.4 points on AlpacaEval 2 and by up to 7.5 points on Arena-Hard. Our top-performing model, built on Llama3-8B-Instruct, achieves a remarkable 53.7 length-controlled win rate on AlpacaEval 2—surpassing Claude 3 Opus on the leaderboard, and a 36.5 win rate on Arena-Hard—making it the strongest 8B open-source model. [^0] Code and models can be found at https://github.com/princeton-nlp/SimPO.
@@ 1 Introduction
Figure 1: SimPO and DPO mainly differ in their reward formulation, as indicated in the shaded box. SimPO outperforms DPO across a wide range of settings on AlpacaEval 2 and Arena-Hard.
@@
Llama3-Instruct-8B-SimPO-v0.2 53.7 47.5 1777
GPT-4 Turbo (11/06) 50.0 50.0 2049
Llama3-Instruct-8B-SimPO 44.7 40.5 1825
@@
Llama3-Instruct-8B-DPO 40.3 37.9 1837
Llama3-Instruct-70B 34.4 33.2 1919
Llama3-Instruct-8B 26.0 25.3 1899
GPT-3.5 Turbo (06/13) 22.7 14.1 1328
@@
  • Simplicity: SimPO does not require a reference model, making it more lightweight and easier to implement compared to DPO and other reference-dependent approaches.
  • Significant performance advantage: Despite its simplicity, SimPO significantly outperforms DPO and its latest variants (e.g., a recent reference-free objective ORPO [40]). The performance advantage is consistent across various training setups and extensive instruction-following benchmarks, including AlpacaEval 2 [53, 26] and the challenging Arena-Hard [52] benchmark. It achieves up to a 6.4 point improvement on AlpacaEval 2 and a 7.5 point improvement on Arena-Hard.
@@
Extensive analysis shows that SimPO utilizes preference data more effectively, leading to a more accurate likelihood ranking of winning and losing responses on a held-out validation set, which in turn translates to a better policy model. As shown in Table 1, we produce a top-performing model, built on top of Llama3-8B-instruct, that achieves a notable 44.7 length-controlled win rate on AlpacaEval 2, outperforming Claude 3 Opus on the leaderboard, and a 33.6 win rate on Arena-Hard, making it the strongest 8B open-source model to date. By training on a preference optimization dataset annotated by a stronger reward model (i.e., ArmoRM [80]), we further produce a stronger v0.2 model with an AlpacaEval 2 LC win rate of 53.7 and an Arena-Hard win rate of 36.5 (Appendix G).
@@ 2.1 Background: Direct Preference Optimization (DPO)
DPO [64] is one of the most popular offline preference optimization methods. Instead of learning an explicit reward model [60], DPO reparameterizes the reward function $r$ using a closed-form expression with the optimal policy:
@@ 2.2 A Simple Reference-Free Reward Aligned with Generation
Using Eq. (1) as the implicit reward expression have the following drawbacks: (1) the requirement of a reference model $\pi_{\text{ref}}$ during training incurs additional memory and computational costs, and (2) there is a discrepancy between the reward being optimized during training and the generation metric used for inference. Specifically, during generation, the policy model $\pi_{\theta}$ is used to generate a sequence that approximately maximizes the average log likelihood, defined as follows:
$p_{\theta}(y\mid x)=\frac{1}{|y|}\log\pi_{\theta}(y\mid x)=\frac{1}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i}).$ (3)
Direct maximization of this metric during decoding is intractable, and various decoding strategies can be used to approximate it, such as greedy decoding [32], beam search [33, 51], nucleus sampling [37], and top-$k$ sampling [28, 38, 63]. Additionally, this metric is commonly used for ranking options in multiple-choice tasks within language models [12, 39, 60]. In DPO, for any triple $(x,y_{w},y_{l})$, satisfying the reward ranking $r(x,y_{w})>r(x,y_{l})$ does not necessarily mean that the likelihood ranking $p_{\theta}(y_{w}\mid x)>p_{\theta}(y_{l}\mid x)$ is met. In fact, only roughly $50\%$ of the triples from the training set satisfy this condition when trained with DPO (see Figure 4(b)).
@@
Naturally, we consider replacing the reward formulation in DPO with $p_{\theta}$ in Eq. (3), so that it aligns with the likehood metric that guides generation. This results in a length-normalized reward:
$r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)=\frac{\beta}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i}),$ (4)
@@ 2.3 The SimPO Objective
In summary, SimPO employs an implicit reward formulation that directly aligns with the generation metric, eliminating the need for a reference model. Additionally, it introduces a target reward margin $\gamma$ to help separating the winning and losing responses. In Appendix E, we provide a gradient analysis of SimPO and DPO to further understand the differences between the two methods.
@@ 3 Experimental Setup
We perform preference optimization with two families of models, Llama3-8B [2] and Mistral-7B [42] under two setups: Base and Instruct.
@@
For the Instruct setup, we use off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct, or mistralai/Mistral-7B-Instruct-v0.2) as the SFT models. [^2] It is unclear whether the released instruct checkpoints have undergone supervised fine-tuning (SFT) or the complete reinforcement learning from human feedback (RLHF) pipeline. For simplicity, we refer to these checkpoints as SFT models. These models have undergone extensive instruction-tuning processes, making them more powerful and robust than the SFT models in the Base setup. However, they are also more opaque because their RLHF procedure is not publicly disclosed. To mitigate the distribution shift between SFT models and the preference optimization process, we generate the preference dataset using the SFT models following [76]. This makes our Instruct setup closer to an on-policy setting. Specifically, we use prompts from the UltraFeedback dataset and regenerate the chosen and rejected response pairs $(y_{w},y_{l})$ with the SFT models. For each prompt $x$, we generate 5 responses using the SFT model with a sampling temperature of 0.8. We then use llm-blender/PairRM [43] to score the 5 responses, selecting the highest-scoring one as $y_{w}$ and the lowest-scoring one as $y_{l}$. For simplicity, we generate data in a single pass instead of iteratively as in [76]. In addition, we create a v0.2 Llama3-Instruct setting by using RLHFlow/ArmoRM-Llama3-8B-v0.1 [80] as the reward model to rank generated data, which yields significantly improved performance (more details in Appendix G).
In summary, we have four setups: Llama3-Base, Llama3-Instruct, Mistral-Base, and Mistral-Instruct. We believe these configurations represent the state-of-the-art, placing our models among the top performers on various leaderboards. We encourage future research to adopt these settings for better and fairer comparisons of different algorithms.
Additionally, we find that tuning hyperparameters is crucial for achieving optimal performance with all the offline preference optimization algorithms, including SimPO. Generally, for SimPO, setting $\beta$ between 2.0 and 2.5 and $\gamma$ between 0.5 and 1.5 leads to good performance across all setups. For more details, please refer to Appendix A.
@@
We primarily assess our models using three of the most popular open-ended instruction-following benchmarks: MT-Bench [94], AlpacaEval 2 [53], [^3] https://tatsu-lab.github.io/alpaca_eval/ and Arena-Hard v0.1 [52]. These benchmarks evaluate the models’ versatile conversational abilities across a diverse set of queries and have been widely adopted by the community (details in Table 2).
AlpacaEval 2 consists of 805 questions from 5 datasets, and MT-Bench covers 8 categories with 80 questions. The most recently released Arena-Hard is an enhanced version of an MT-Bench, incorporating 500 well-defined technical problem-solving queries. We report scores following each benchmark’s evaluation protocol. For AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC) [26]. The LC metric is specifically designed to be robust against model verbosity. For Arena-Hard, we report the win rate (WR) against the baseline model. For MT-Bench, we report the average MT-Bench score with GPT-4 and GPT-4-Preview-1106 as the judge model. [^4] GPT-4-Preview-1106 produces more accurate reference answers and judgments compared to GPT-4. For decoding details, please refer to Appendix A. We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks [9], with additional details in in Appendix B.
@@
R-DPO [62] $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
@@
We compare SimPO with other offline preference optimization methods listed in Table 3.555Many recent studies [86, 72] have extensively compared DPO and PPO [68]. We will leave the comparison of PPO and SimPO to future work. RRHF [87] and SLiC-HF [92] are ranking losses. RRHF uses length-normalized log-likelihood, similar to SimPO’s reward function, while SLiC-HF uses log-likelihood directly and includes an SFT objective. IPO [6] is a theoretically grounded approach method that avoids DPO’s assumption that pairwise preferences can be replaced with pointwise rewards. CPO [84] uses sequence likelihood as a reward and trains alongside an SFT objective. KTO [27] learns from non-paired preference data. ORPO [40] [^6] ORPO can directly train on preference data without the SFT stage. For fair comparisons, we start ORPO from the same SFT checkpoints as other baselines, which should yield better results than starting from a base checkpoint. introduces a reference-model-free odd ratio term to directly contrast winning and losing responses with the policy model and jointly trains with the SFT objective. R-DPO [62] is a modified version of DPO that includes an additional regularization term to prevent exploitation of length. We thoroughly tune the hyperparameters for each baseline and report best performance. We find that many variants of DPO do not empirically present an advantage over standard DPO. Further details can be found in Appendix A.
@@ 4.1 Main Results and Ablations
Method Llama3-Base (8B) Llama3-Instruct (8B)
@@
In Table 5, we demonstrate results from ablating each key design of SimPO: (1) removing length normalization in Eq. (4) (i.e., w/o LN); (2) setting the target reward margin to be 0 in Eq. (6) (i.e., $\gamma=0$). Removing the length normalization has the most negative impact on the results. Our examination reveals that this leads to the generation of long and repetitive patterns, substantially degrading the overall quality of the output (See Appendix D). Setting $\gamma$ to 0 yields also leads to a performance degradation compared to SimPO, indicating that it is not the optimal target reward margin. In the following subsections, we conduct in-depth analyses to better understand both design choices.
@@ 4.3 The Impact of Target Reward Margin in SimPO
We investigate how the target reward margin $\gamma$ in SimPO affects the reward accuracy on a held-out set and win rate on AlpacaEval 2, presenting the results in Figure 3(a). Reward accuracy is measured as the percentage of preference pairs where the winning response ends up having a higher reward for the wining response than the losing response (i.e., $r(x,y_{w})>r(x,y_{l})$). We observe that reward accuracy increases with $\gamma$ on both benchmarks, indicating that enforcing a larger target reward margin effectively improves reward accuracy. However, the win rate on AlpacaEval 2 first increases and then decreases with $\gamma$, suggesting that generation quality is not solely determined by the reward margin.
@@
Figure 5: Comparison between SimPO and DPO (continued). (a) With different $\beta$ in DPO and SimPO, KL divergence from the policy model to the reference model on winning responses. (b) AlpacaEval2 LC win rate of DPO and SimPO with different $\beta$. (c) Runtime and memory usage for DPO and SimPO.
@@
Although the DPO reward expression $r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ (with the partition function excluded) lacks an explicit term for length normalization, the logarithmic ratio between the policy model and the reference model can serve to implicitly counteract length bias. As shown in Table 6 and Figure 4(a), employing DPO reduces the Spearman correlation coefficient between average log likelihood and response length compared to the approach without any length normalization (referred to as “SimPO w/o LN”). However, it still exhibits a stronger positive correlation when compared to SimPO. [^8] Note that this correlation does not fully reflect the generation length. Despite DPO showing a stronger correlation, the length of its generated responses is comparable to or even slightly shorter than those of the SimPO models. Please find more details in Appendix D.
@@
In Figure 4(c), we compare the reward accuracy of SimPO and DPO, assessing how well their final learned rewards align with preference labels on a held-out set. We observe that SimPO consistently achieves higher reward accuracy than DPO, suggesting that our reward design facilitates more effective generalization and leads to higher quality generations.
@@
In Figure 5(a), we present the KL divergence between the policy model trained with DPO and SimPO and the reference model with different $\beta$, measured on the winning responses from a held-out set during training. Figure 5(b) shows the corresponding AlpacaEval 2 LC win rate. Although SimPO does not apply any form of regularization against the reference model, the KL divergence of SimPO is reasonably small. Increasing $\beta$ reduces the KL divergence for both DPO and SimPO, with DPO exhibiting a more pronounced reduction at higher $\beta$ values. In this particular setting (Mistral-base), Figure 5(b) demonstrates that a smaller $\beta$ can yield better AlpacaEval 2 performance, despite the higher KL divergence. [^9] We observe that in some settings (e.g., Llama3-Instruct), having a large $\beta$ (e.g., $\beta=10$) leads to better performance. We hypothesize that when the reference model is weak, strictly constraining the policy model to the reference model may not be beneficial. As a caveat, while we did not observe any training collapse or degeneration with proper tuning, in principle, training with SimPO could potentially lead to reward hacking since SimPO does not regularize against the reference model. In such a scenario, the model might achieve a low loss but degenerate in sequence generation.
@@
Another benefit of SimPO is its efficiency as it does not use a reference model. Figure 5(c) illustrates the overall run time and per-GPU peak memory usage of SimPO and DPO in the Llama3-Base setting using 8×H100 GPUs. Compared to a vanilla DPO implementation, [^10] DPO can be as memory efficient as SimPO if it were implemented to separate the forward passes of the reference model from the actual preference optimization. However, this implementation is not standard practice. SimPO cuts run time by roughly 20% and reduces GPU memory usage by about 10%, thanks to eliminating forward passes with the reference model.

### v3

@@ Abstract
Direct Preference Optimization (DPO) is a widely used offline preference optimization algorithm that reparameterizes reward functions in reinforcement learning from human feedback (RLHF) to enhance simplicity and training stability. In this work, we propose SimPO, a simpler yet more effective approach. The effectiveness of SimPO is attributed to a key design: using the average log probability of a sequence as the implicit reward. This reward formulation better aligns with model generation and eliminates the need for a reference model, making it more compute and memory efficient. Additionally, we introduce a target reward margin to the Bradley-Terry objective to encourage a larger margin between the winning and losing responses, further improving the algorithm’s performance. We compare SimPO to DPO and its recent variants across various state-of-the-art training setups, including both base and instruction-tuned models such as Mistral, Llama 3, and Gemma 2. We evaluate on extensive chat-based evaluation benchmarks, including AlpacaEval 2, MT-Bench, and Arena-Hard. Our results demonstrate that SimPO consistently and significantly outperforms existing approaches without substantially increasing response length. Specifically, SimPO outperforms DPO by up to 6.4 points on AlpacaEval 2 and by up to 7.5 points on Arena-Hard. Our top-performing model, built on Gemma-2-9B-it, achieves a 72.4% length-controlled win rate on AlpacaEval 2, a 59.1% win rate on Arena-Hard, and ranks 1st on Chatbot Arena among $<$10B models with real user votes. [^0] Code and models can be found at https://github.com/princeton-nlp/SimPO.
@@ 1 Introduction
Figure 1: SimPO and DPO mainly differ in their reward formulation, as indicated in the shaded box. SimPO outperforms DPO significantly across a range of settings on AlpacaEval 2 and Arena-Hard.
@@
Gemma-2-9B-it 51.1 38.1 1571
Llama-3-8B-Instruct-SimPO 44.7 40.5 1825
@@
Llama-3-8B-Instruct-DPO 40.3 37.9 1837
Llama-3-70B-Instruct 34.4 33.2 1919
Llama-3-8B-Instruct 26.0 25.3 1899
@@
  • Simplicity: SimPO does not require a reference model, making it more lightweight and easier to implement compared to DPO and other reference-based methods.
  • Significant performance advantage: Despite its simplicity, SimPO significantly outperforms DPO and its latest variants (e.g., a recent reference-free objective ORPO [42]). The performance advantage is consistent across various training setups and extensive chat-based evaluations, including AlpacaEval 2 [55, 28] and the challenging Arena-Hard [54] benchmark. It achieves up to a 6.4 point improvement on AlpacaEval 2 and a 7.5 point improvement on Arena-Hard compared to DPO (Figure 1).
@@
Extensive analysis shows that SimPO utilizes preference data more effectively, leading to a more accurate likelihood ranking of winning and losing responses on a held-out validation set, which in turn translates to a better policy model. As shown in Table 1, our Gemma-2-9B-it-SimPO model achieves state-of-the-art performance, with a $72.4\%$ length-controlled win rate on AlpacaEval 2 and a $59.1\%$ win rate on Arena-Hard, establishing it as the strongest open-source model under 10B parameters. Most notably, when evaluated on Chatbot Arena [17] with real user votes, our model significantly improved upon the initial Gemma-2-9B-it model, advancing from 36th to 25th place and ranking first among all $<$10B models on the leaderboard. [^1] As of September 16th, 2024.
@@ 2.1 Background: Direct Preference Optimization (DPO)
DPO [66] is one of the most popular preference optimization methods. Instead of learning an explicit reward model [62], DPO reparameterizes the reward function $r$ using a closed-form expression with the optimal policy:
@@ 2.2 A Simple Reference-Free Reward Aligned with Generation
Using Eq. (1) as the implicit reward has the following drawbacks: (1) it requires a reference model $\pi_{\text{ref}}$ during training, which incurs additional memory and computational costs; and (2) it creates a mismatch between the reward optimized in training and the log-likelihood optimized during inference, where no reference model is involved. This means that in DPO, for any triple $(x,y_{w},y_{l})$, satisfying the reward ranking $r(x,y_{w})>r(x,y_{l})$ does not necessarily mean that the likelihood ranking $p_{\theta}(y_{w}\mid x)>p_{\theta}(y_{l}\mid x)$ is met (here $p_{\theta}$ is the average log-likelihood in Eq. (3)). In our experiments, we observed that only $\sim 50\%$ of the triples from the training set satisfy this condition when trained with DPO (Figure 4(b)). This observation aligns with a concurrent work [14], which finds that existing models trained with DPO exhibit random ranking accuracy in terms of average log-likelihood, even after extensive preference optimization.
@@
One solution is to use the summed token log probability as the reward, but this suffers from length bias–longer sequences tend to have lower log probabilities. Consequently, when $y_{w}$ is longer than $y_{l}$, optimizing the summed log probability as a reward forces the model to artificially inflate probabilities for longer sequences to ensure $y_{w}$ receives a higher reward than $y_{l}$. This overcompensation increases the risk of degeneration. To address this issue, we consider using the average log-likelihood as the implicit reward:
$p_{\theta}(y\mid x)=\frac{1}{|y|}\log\pi_{\theta}(y\mid x)=\frac{1}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i}).\vspace{-0.5em}$ (3)
This metric is commonly used for ranking options in beam search [35, 53] and multiple-choice tasks within language models [12, 41, 62]. Naturally, we consider replacing the reward formulation in DPO with $p_{\theta}$ in Eq. (3), so that it aligns with the likelihood metric that guides generation. This results in a length-normalized reward:
$r_{\text{SimPO}}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)=\frac{\beta}{|y|}\sum_{i=1}^{|y|}\log\pi_{\theta}(y_{i}\mid x,y_{<i}),\vspace{-0.1em}$ (4)
@@ 2.3 The SimPO Objective
In summary, SimPO employs an implicit reward formulation that directly aligns with the generation metric, eliminating the need for a reference model. Additionally, it introduces a target reward margin $\gamma$ to help separating the winning and losing responses. In Appendix F, we provide a gradient analysis of SimPO and DPO to further understand the differences between the two methods.
Preventing catastrophic forgetting without KL regularization.
Although SimPO does not impose KL regularization, we find that a combination of practical factors ensures effective learning from preference data while maintaining generalization, leading to an empirically low KL divergence from the reference model. These factors are: (1) a small learning rate, (2) a preference dataset that covers diverse domains and tasks, and (3) the intrinsic robustness of LLMs to learn from new data without forgetting prior knowledge. We present KL divergence experiments in Section 4.4.
@@ 3 Experimental Setup
We perform preference optimization with two families of models, Llama-3-8B [2] and Mistral-7B [44], under two setups: Base and Instruct. In this section, our goal is to understand the performance of SimPO vs. other preference optimization methods in different experimental setups. Our strongest model is based on Gemma-2-9B (Instruct setup) with a stronger reward model, RLHFlow/ArmoRM-Llama3-8B-v0.1 [84] (Table 1). We will present and discuss these results in Appendix J.
@@
For the Instruct setup, we use an off-the-shelf instruction-tuned model (i.e., meta-llama/Meta-Llama-3-8B-Instruct, or mistralai/Mistral-7B-Instruct-v0.2) as the SFT models. [^3] It is unclear whether the released instruct checkpoints have undergone supervised fine-tuning (SFT) or the complete RLHF pipeline. For simplicity, we refer to these checkpoints as SFT models. These models have undergone extensive instruction-tuning processes, making them more powerful and robust than the SFT models in the Base setup. However, they are also more opaque because their RLHF procedure is not publicly disclosed. To mitigate the distribution shift between SFT models and the preference optimization process, we generate the preference dataset using the SFT models following [79]. This makes our Instruct setup closer to an on-policy setting. Specifically, we use prompts from the UltraFeedback dataset and regenerate the chosen and rejected response pairs $(y_{w},y_{l})$ with the SFT models. For each prompt $x$, we generate 5 responses using the SFT model with a sampling temperature of 0.8. We then use llm-blender/PairRM [45] to score the 5 responses, selecting the highest-scoring one as $y_{w}$ and the lowest-scoring one as $y_{l}$. We only generated data in a single pass instead of iteratively as in [79].444 We also experimented with using a stronger reward model, RLHFlow/ArmoRM-Llama3-8B-v0.1 [84], to rank generated data, which yields significantly improved performance (see Appendix H and Appendix J). This is the reward model we used in our Gemma 2 experiments.
In summary, we have four setups: Llama-3-Base, Llama-3-Instruct, Mistral-Base, and Mistral-Instruct. We believe these configurations represent the state-of-the-art, placing our models among the top performers on various leaderboards. We encourage future research to adopt these settings for better and fairer comparisons of different algorithms. Additionally, we find that tuning hyperparameters is crucial for achieving optimal performance with all the offline preference optimization algorithms, including DPO and SimPO. Generally, for SimPO, setting $\beta$ between 2.0 and 2.5 and $\gamma$ between 0.5 and 1.5 leads to good performance across all setups. For more details, please refer to Appendix B.
@@
We primarily assess our models using three of the most popular open-ended instruction-following benchmarks: MT-Bench [99], AlpacaEval 2 [55], and Arena-Hard v0.1 [54]. These benchmarks evaluate the models’ versatile conversational abilities across a diverse set of queries and have been widely adopted by the community (details in Table 2). AlpacaEval 2 consists of 805 questions from 5 datasets, and MT-Bench covers 8 categories with 80 questions. The most recently released Arena-Hard is an enhanced version of an MT-Bench, incorporating 500 well-defined technical problem-solving queries. We report scores following each benchmark’s evaluation protocol. For AlpacaEval 2, we report both the raw win rate (WR) and the length-controlled win rate (LC) [28]. The LC metric is specifically designed to be robust against model verbosity. For Arena-Hard, we report the win rate (WR) against the baseline model. For MT-Bench, we report the average MT-Bench score with GPT-4 and GPT-4-Preview-1106 as the judge model. [^5] GPT-4-Preview-1106 produces more accurate reference answers and judgments compared to GPT-4. For decoding details, please refer to Appendix B. We also evaluate on downstream tasks from the Huggingface Open Leaderboard benchmarks [9], with additional details in in Appendix C.
@@
R-DPO [64] $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}+\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
@@
We compare SimPO with other offline preference optimization methods listed in Table 3.666Many recent studies [90, 74] have extensively compared DPO and PPO [70]. We will leave the comparison of PPO and SimPO to future work. RRHF [91] and SLiC-HF [96] are ranking losses. RRHF uses length-normalized log-likelihood, similar to SimPO’s reward function, while SLiC-HF uses log-likelihood directly and includes an SFT objective. IPO [6] is a theoretically grounded approach method that avoids DPO’s assumption that pairwise preferences can be replaced with pointwise rewards. CPO [88] uses sequence likelihood as a reward and trains alongside an SFT objective. KTO [29] learns from non-paired preference data. ORPO [42] [^7] ORPO can directly train on preference data without the SFT stage. For fair comparisons, we start ORPO from the same SFT checkpoints as other baselines, which yields better results than starting from base checkpoints. introduces a reference-model-free odd ratio term to directly contrast winning and losing responses with the policy model and jointly trains with the SFT objective. R-DPO [64] is a modified version of DPO that includes an additional regularization term to prevent exploitation of length. We thoroughly tune the hyperparameters for each baseline and report the best performance. We find that many variants of DPO do not empirically present an advantage over standard DPO. Further details can be found in Appendix B.
@@ 4.1 Main Results and Ablations
Method Llama-3-Base (8B) Llama-3-Instruct (8B)
@@
In Table 5, we demonstrate results from ablating each key design of SimPO: (1) removing length normalization in Eq. (4) (i.e., w/o LN); (2) setting the target reward margin to be 0 in Eq. (6) (i.e., $\gamma=0$). Removing the length normalization has the most negative impact on the results. Our examination reveals that this leads to the generation of long and repetitive patterns, substantially degrading the overall quality of the output (See Appendix E). Setting $\gamma$ to 0 yields also leads to a performance degradation compared to SimPO, indicating that it is not the optimal target reward margin. In the following subsections, we conduct in-depth analyses to better understand both design choices.
@@ 4.3 The Impact of Target Reward Margin in SimPO
We investigate how the target reward margin $\gamma$ in SimPO affects the reward accuracy on a held-out set and win rate on AlpacaEval 2, presenting the results in Figure 3(a). Reward accuracy is measured as the percentage of preference pairs where the winning response ends up having a higher reward for the winning response than the losing response (i.e., $r(x,y_{w})>r(x,y_{l})$). We observe that reward accuracy increases with $\gamma$ on both benchmarks, indicating that enforcing a larger target reward margin effectively improves reward accuracy. However, the win rate on AlpacaEval 2 first increases and then decreases with $\gamma$, suggesting that generation quality is not solely determined by the reward margin.
@@
Figure 5: Comparison between SimPO and DPO (continued). (a) With different $\beta$ in DPO and SimPO, KL divergence from the policy model to the reference model on $y_{w}$. (b) AlpacaEval2 LC win rate of DPO and SimPO with different $\beta$. (c) Runtime and memory usage for DPO and SimPO.
@@
Although the DPO reward expression $r(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ (with the partition function excluded) lacks an explicit term for length normalization, the logarithmic ratio between the policy model and the reference model can serve to implicitly counteract length bias. As shown in Table 6 and Figure 4(a), employing DPO reduces the Spearman correlation coefficient between average log likelihood and response length compared to the approach without any length normalization (referred to as “SimPO w/o LN”). However, it still exhibits a stronger positive correlation when compared to SimPO. [^9] Note that this correlation does not fully reflect the generation length. Despite DPO showing a stronger correlation, the length of its generated responses is comparable to or even slightly shorter than those of the SimPO models. Please find more details in Appendix E.
@@
In Figure 4(c), we compare the reward accuracy of SimPO and DPO, assessing how well their final learned rewards align with preference labels on a held-out set. SimPO consistently achieves higher reward accuracy than DPO, suggesting that our reward design facilitates better generalization and leads to higher quality generations.
@@
In Figure 5(a), we present the KL divergence between the policy model trained with DPO and SimPO and the reference model with different $\beta$, measured on the winning responses from a held-out set during training. Figure 5(b) shows the corresponding AlpacaEval 2 LC win rate. Although SimPO does not apply any form of regularization against the reference model, the KL divergence of SimPO is reasonably small. Increasing $\beta$ reduces the KL divergence for both DPO and SimPO, with DPO exhibiting a more pronounced reduction at higher $\beta$ values. In this particular setting (Mistral-base), Figure 5(b) demonstrates that a smaller $\beta$ can improve AlpacaEval 2 performance, despite the higher KL divergence. [^10] We observe that in some settings (e.g., Llama-3-Instruct), a large $\beta$ (e.g., $\beta=10$) leads to better performance. We hypothesize that when the reference model is weak, strictly constraining the policy model to the reference model may not be beneficial. As a caveat, while we did not observe any training collapse or degeneration with proper tuning, in principle, SimPO could potentially lead to reward hacking without explicit regularization against the reference model. In such a scenario, the model might achieve a low loss but degenerate.
@@
Another benefit of SimPO is its efficiency as it does not use a reference model. Figure 5(c) illustrates the overall run time and per-GPU peak memory usage of SimPO and DPO in the Llama-3-Base setting using 8×H100 GPUs. Compared to a vanilla DPO implementation, [^11] DPO can be as memory efficient as SimPO if it were implemented to separate the forward passes of the reference model from the actual preference optimization. However, this implementation is not standard practice. SimPO cuts run time by roughly 20% and reduces GPU memory usage by about 10%, thanks to eliminating forward passes with the reference model.